Two dice are thrown at a time. The probability of the difference of the numbers shown on the dice is 1 is

Grade:9

1 Answers

Arun
6 years ago
Dear Devansh
 
Total favourable outcomes = { (1,2),(2,1),(3,2)(2,3)(3,4)(4,3)(4,5)(5,4)(5,6)(6,5)}
 
Hence no of total outcomes = 10
 
Then probability = 10/36 = 5/18
 
 
Regards
Arun (askIITians forum expert)